OpenStudy (anonymous):

Write the complex number in the form a + bi. 3(cos 60° + i sin 60°)

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@aum

myininaya (myininaya):

Use the unit circle to evaluate cos(60) and sin(60)

OpenStudy (aum):

\[ 3(\cos 60° + i \sin 60°) = 3\left (\frac{\sqrt{3}}{2} + i\frac 12 \right ) = \frac{3\sqrt{3}}{2} + i\frac {3}{2} \]

OpenStudy (anonymous):

i only see question mark symbols next to the numbers. did you type anything besides the question?

OpenStudy (aum):

Refresh the page.

OpenStudy (aum):

\[ 3(\cos 60 + i \sin 60) = 3\left (\frac{\sqrt{3}}{2} + i\frac 12 \right ) = \frac{3\sqrt{3}}{2} + i\frac {3}{2} \]

OpenStudy (anonymous):

oh i see so you just plugged in the unit circle values?

OpenStudy (aum):

yes.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

do you think you could help me with another one